&12.2.15-T Question Help A study reported that 48% of men age 62 to 64 were still working, 34% of women age 62 to 64 were stiHow are the last two calculated?
Answer #1

Applying chi square test:

Oi 62-64 men 62-64 women 65-69 men 65-69 women Total
working 48 34 27 23 132
not working 52 66 73 77 268
total 100 100 100 100 400
Expected Ei=row total*column total/grand total 62-64 men 62-64 women 65-69 men 65-69 women Total
working 33.00 33.00 33.00 33.00 132
not working 67.00 67.00 67.00 67.00 268
total 100 100 100 100 400
chi square    χ2 =(Oi-Ei)2/Ei 62-64 men 62-64 women 65-69 men 65-69 women Total
working 6.8182 0.0303 1.0909 3.0303 10.970
not working 3.3582 0.0149 0.5373 1.4925 5.403
total 10.176 0.045 1.628 4.523 16.373

from above chi square test statistic X2 = 16.373

degree of freedom(df) =(rows-1)*(columns-1)= 3

for 3 df and above test statistic ; p value =0.001